# eprint Nisho-2-2025

### v1

@@ Abstract
Obviously, the method of the axion detection using cylinder sample with small electric conductivity is very powerful. This is because the axion–electromagnetic coupling induces an oscillating electric current throughout the entire cylindrical sample, producing sufficiently large dissipation power $P$ for detection.
@@
Summarized, we have proposed a method for axion detection using a cylindrical sample with choosing appropriate low conductivity $(\sigma=10^{-3}\text{-}10^{-2}\text{eV})$ sensitive to microwave signals in the frequency range $m_{a}/2\pi=24\text{-}240\text{GHz}$. For such conductivity, the axion-induced electric current penetrates the bulk, yielding substantial energy dissipation and a high signal-to-noise ratio, demonstrating the feasibility of detecting dark matter axions under realistic experimental conditions. The method is also effective for the search of dark photon.

### v2

@@ Abstract
We find that the method of the axion detection using cylinder sample with small electrical conductivity is very powerful. This is because the axion–electromagnetic coupling induces an oscillating electric field $|E_{a}|/2\epsilon$ throughout the entire cylindrical sample, producing sufficiently large electric current $I\propto\sigma(|E_{a}|/\epsilon)^{2}R^{2}$ for the detection. Furthermore, the sample with small electrical conductivity has small thermal noise compared with the one with large electrical conductivity.
@@
Summarized, we have proposed a method for axion detection with a cylindrical sample choosing appropriate small conductivity $(\sigma=10^{-3}\text{-}10^{-2}\text{eV})$ sensitive to microwave signals in the frequency range $m_{a}/2\pi=24\text{-}240\text{GHz}$. With such conductivity, the axion induces large electric current flowing the bulk of the cylinder and causes a high signal-to-noise ratio of the current. Therefore, the method demonstrates the feasibility of detecting dark matter axions under realistic experimental conditions. It is also effective for the search of dark photon.